OpenStudy (anonymous):

Find the variation constant and an equation of variation where y varies directly as x and y = 27 when x = 3

OpenStudy (shaik0124):

k=27/3=9

OpenStudy (shaik0124):

sincey=kx

OpenStudy (shaik0124):

y=kx,k=27/3=9

OpenStudy (anonymous):

what does y =?

OpenStudy (anonymous):

y=9x